OpenStudy (anonymous):

Find an anti-derivative for each of the following -2/x^3

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\Large \int\limits\limits_{}^{}\frac{-2}{x^3} dx\] sorry, I had to fix my notation

OpenStudy (anonymous):

Well first you gotta factor out the -2 the integrand

OpenStudy (anonymous):

Extricate it out?

OpenStudy (anonymous):

Factor, but if you want to say that lol, it must go on the outside though

OpenStudy (primeralph):

"Extricate it out" ? Dude, calm down.

OpenStudy (anonymous):

\[\Large -2\int\limits_{}^{}\frac{1}{x^3} dx\]

OpenStudy (anonymous):

ok*

OpenStudy (anonymous):

why you put 1 ?

OpenStudy (anonymous):

Because I factored the -2 out, since I took it out, you have to leave a 1, its like you multiply it in there -2*1=-2 its the same thing.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

anyways you have to change the fraction to an x with a negative exponent to use the Power Rule for integration

OpenStudy (anonymous):

what's the Power Rule for integration ?

OpenStudy (anonymous):

Do you know the power rule for integration? \[\Huge \int\limits_{}^{}x^{n}~dx=\frac{x^{n+1}}{n+1}\]

OpenStudy (anonymous):

Because the first time I am studying integration

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

But we have no shrine and the extension .. how we ride it rule?

OpenStudy (anonymous):

Alright now remember we have \[\Large -2\int\limits\limits_{}^{}\frac{1}{x^3} dx \] We have to change the \[\Large \frac{1}{x^3} ~to~x^{-3}\] \[\Large -2\int\limits\limits\limits_{}^{}{x^{-3}} dx\] \[\Large -2(\frac{x^{-3+1}}{-3+1}) +C\]

OpenStudy (anonymous):

aha thanks a lot

OpenStudy (anonymous):

Don't forget you gotta fix it up \[\Large -2\frac{x^{-2}}{-2} +C\] \[\Large x^{-2}+C \] \[\Large \frac{1}{x^2}+C\] Do you know why you have to put a C

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Alright, good luck, it only gets harder!!!